OpenStudy (lilsis76):

convert the given polar coordinates to rectangular coordinates.(10, pi/5) i know x=rcosQ y=rsinQ i guess i plug in 10 cos(pi/5)? right??

OpenStudy (anikay):

oh god why >.>

OpenStudy (lilsis76):

HAHA

OpenStudy (lilsis76):

Im sorry if i am completely blank

OpenStudy (anikay):

this stuff was the worst part of my junior year xD

OpenStudy (lilsis76):

HAHA, dang how old are you? i never took this in HighSchool. my highest was Algebra

OpenStudy (anikay):

I'm 17. and yes, to get x you plug in \[10\cos (\pi/5)\] and for y \[10\sin (\pi/5)\]

OpenStudy (lilsis76):

well i tried that x=10 cos pi/5 --> 9.99 ? y=10 sin pi/5--> .0349

OpenStudy (anikay):

yup I think so

OpenStudy (lilsis76):

okay, i have a question tho now. i got this question wrong and the instructor circled my pi/5 and wrote radians. isnt it in radians?

OpenStudy (anikay):

yes, should it be degress?

OpenStudy (lilsis76):

lol it doesnt say, but i dont know why she circledit lol. THANK YOU OH SO VERY MUCH!!!!! i really appreciate it. UGH i gotta do better on this test make up.

OpenStudy (lilsis76):

Have a goodnite!

OpenStudy (anikay):

i got a test make-up friday and a test tomorrow

OpenStudy (anikay):

Calc, is much easier imo

OpenStudy (lilsis76):

oh boy, :/ i would def. fail that. lol i hope so i have to get all the way to calc 2 i think. Well im logging off. thank you very much. Have a goodnite

OpenStudy (anikay):

if you ever need calc help, I'm an @ away. night

OpenStudy (lilsis76):

haha

OpenStudy (phi):

*** well i tried that x=10 cos pi/5 --> 9.99 ? y=10 sin pi/5--> .0349 *** right idea, but you had your calculator in "degree mode" rather than "radian mode" so you are getting the wrong answer the safest way to do this is change pi/5 radians to degrees you do this by multiplying by 180/pi \[ \frac{\cancel{\pi}}{5} \text{ rads} \cdot \frac{180}{\cancel{\pi}}= \frac{180}{5}= 36º \] now, with your calculator in "DEGREE MODE" do 10*cos(36)= 8.08 10*sin(36)= 5.88 your rectangular coordinates are (8.08, 5.88)

OpenStudy (anikay):

oh right, because theta always gives a polar answer, thanks :D